OpenStudy (sleepyjess):

Find a cubic function with the given zeros. \(\large\sqrt{7},~ -\sqrt{7}, ~-4\)

OpenStudy (campbell_st):

so are the roots \[x = -\sqrt{7}, \sqrt{7}, 4\]

OpenStudy (sleepyjess):

yes

OpenStudy (sleepyjess):

-4 not 4

OpenStudy (campbell_st):

ok... then its simply \[P(x) = (x - \sqrt{7})(x + \sqrt{7})(x - 4)\] you should notice the 1st 2 binomials are the difference of 2 squares hope it helps

OpenStudy (campbell_st):

ok... so if it's x = -4 the last binomial changes to (x + 4) then just multiply them out

OpenStudy (sleepyjess):

(x-\(\sqrt{7}\))(x+\(\sqrt{7}\)) \(x^2-x\sqrt{7}+x\sqrt{7}-\sqrt{49?~(idk)}\)

OpenStudy (sleepyjess):

then multiply that by x+4 right?

OpenStudy (sleepyjess):

is the last part of my equation right so far?

OpenStudy (campbell_st):

well yes the 1st and 2 binomials expand to \[P(x) = (x^2 - 7)(x + 4) \]

OpenStudy (sleepyjess):

Ok i was just making sure with everything in it first

OpenStudy (sleepyjess):

(\(x^2-7)(x+4)\) \(x^3+4x^2-7x-28\)

OpenStudy (campbell_st):

that's great... don't for get to let it equal something f(x) or something like that

OpenStudy (sleepyjess):

ok thank you so much